OpenStudy (anonymous):

Simplify the expression 5^3 × 5^8.

OpenStudy (anonymous):

Do you add the exponents or multiply them?

jigglypuff314 (jigglypuff314):

5^3 is 5*5*5 where * represents multiplication

OpenStudy (anonymous):

So you add the exponents, not multiply?

jigglypuff314 (jigglypuff314):

yes :) x^3 * x^5 = x^(3+5) = x^8

OpenStudy (anonymous):

so it'd be 5^11?

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (anonymous):

Thanks so much!

jigglypuff314 (jigglypuff314):

you're welcome :)